How many solutions does this system have?
–3x + 6y = 10
–3x + 6y = –4

Found 2 solutions by flame8855, ewatrrr:
Answer by flame8855(424)

3x - 6y = -10
–3x + 6y = –4
0 = -14
no solution

Answer by ewatrrr(24785)

 
Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
How many solutions does this system have?
–3x + 6y = 10 OR y = .5x + 10/6
–3x + 6y = –4 OR y = .5x + -4/6
Note: slopes are Indentical, Lines are parallel, NO solutions for the system.